# Bright single-photon source in a silicon chip by nanoscale positioning of a color center in a microcavity

### v1

@@ Conclusion
In conclusion, we have demonstrated a source of near-infrared linearly-polarized single photons obtained by nanoscale positioning of a W center in a SOI CBG cavity. Si+ implantation through a mask patterned with nanoholes allows for selective creation of single W centers, positioned in the vicinity of the main antinode of the resonant mode. The cavity-coupled emitter has a PL intensity at saturation of around 1.2 Mcps together with a value of around 0.06 under above-gap CW excitation, with a ZPL fraction of 98.6. Under weak pumping condition, the triggered emission of single-photon pulses is demonstrated with 93 purity. These results highlight the great potential of W centers for silicon quantum photonics. They pave the way towards the scalable integration of bright sources of indistinguishable single photons on integrated photonic chips, based on single W centers deterministically positioned in high-Fp waveguide-coupled photonic crystal nanocavities.
@@ Methods
The implantation of Si+ ions is performed at an energy of 70 keV, with a dose of 1011 cm-2 and 7° tilt to avoid channeling.
@@
Continuous-wave and pulsed optical excitation are provided by a semiconductor laser (PicoQuant LDH-IB-485-B) operating around 485 nm.
@@
We thank Guillaume Cassabois and Vincent Jacques for fruitful discussions, as well as Krithika V. R. and Alicia Morandini for experimental support. This work is supported by the Plan France 2030 through the project OQuLuS (No. ANR-22-PETQ-0013), the French National Research Agency (ANR) through the projects OCTOPUS (No. ANR-18-CE47-0013-01) and WOUAH (No. ANR-24-CE47-4667), CEA through the PTC-MP “W-TeQ” internal project, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project SILEQS, Grant No. 101042075). B.L. is supported by the “Program QuantForm-UGA No. ANR-21-CMAQ-003 France 2030” and “Laboratoire d’Excellence LANEF No. ANR-10-LABX-51-01”. C.E. is supported by the MSCA Cofund QuanG (Grant Number: 101081458), funded by the European Union. The authors acknowledge the assistance of the staff of the Grenoble Advanced Technological Platform (PTA).
See the Supporting Information for additional details about simulation of ion implantation, optical properties of single W centers in unpatterned SOI, finite-difference time-domain simulations, single W centers hunting strategy, and estimation of the zero-phonon emission rate for the CBG-coupled W center.

### v2

@@ Conclusion
In conclusion, we have demonstrated a source of near-infrared linearly-polarized single photons obtained by nanoscale positioning of a W center in a SOI CBG cavity. Si+ implantation through a mask patterned with nanoholes allows for selective creation of single W centers, positioned in the vicinity of the main antinode of the resonant mode. The cavity-coupled emitter has a PL intensity at saturation of Mcps together, with a value inferior to 0.06 under above-gap CW excitation up to , and with a ZPL fraction of . Under weak pumping condition, the triggered emission of single-photon pulses is demonstrated with purity. These results highlight the great potential of W centers for silicon quantum photonics. They pave the way towards the scalable integration of bright sources of indistinguishable single photons on integrated photonic chips, based on single W centers deterministically positioned in high- waveguide-coupled photonic crystal nanocavities.
@@ Methods
The implantation of Si+ ions is performed at an energy of 70 keV, with a dose of 1011 cm-2 and 7° tilt angle with respect to the normal of the sample to avoid channeling.
@@
Continuous-wave and pulsed optical excitation are provided by a semiconductor laser (PicoQuant LDH-IB-485-B) operating at 485 nm.
@@
Data availability
The data supporting the findings of this study are available from the corresponding author upon reasonable request.
Acknowledgements.
We thank Guillaume Cassabois and Vincent Jacques for fruitful discussions, as well as Krithika V. R. and Alicia Morandini for experimental support. This work is supported by the Plan France 2030 through the project OQuLuS (No. ANR-22-PETQ-0013), the French National Research Agency (ANR) through the projects OCTOPUS (No. ANR-18-CE47-0013-01) and WOUAH (No. ANR-24-CE47-4667), CEA through the PTC-MP “W-TeQ” internal project, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project SILEQS, Grant No. 101042075). B.L. is supported by the “Program QuantForm-UGA No. ANR-21-CMAQ-003 France 2030” and “Laboratoire d’Excellence LANEF No. ANR-10-LABX-51-01”. C.E. is supported by the MSCA Cofund QuanG (Grant Number: 101081458), funded by the European Union. The authors acknowledge the assistance of the staff of the Grenoble Advanced Technological Platform (PTA).Supplementary Information Available
See the Supporting Information for additional details about simulation of ion implantation, optical properties of single W centers in unpatterned SOI, finite-difference time-domain simulations, estimation of the quantum efficiency and zero-phonon emission rate of the W-1 center, SEM image, and photon autocorrelation measurement at high power.